gavinbacon2:

can someone help me (5+5x3-7x4) + (x+3x3-4x4-8)= 4 3

gavinbacon2:

the 4 and the 3 is squared to the power

Angle:

do you mean: \((5+5x^3-7x^4)+(x +3x^3-4x^4 - 8) = 4^3\)

gavinbacon2:

no we are doing polynomial

Angle:

I did write a polynomial.... I am not sure what your question is asking, can you take a screenshot or use the draw tool?

gavinbacon2:

yes

gavinbacon2:

Angle:

OH that makes more sense, thank you

Angle:

In this problem, we want to "combine like terms" the variables that have the same degree value are considered "like terms"

Angle:

so we can combine: \(-7x^4\) with \(-4x^4\) because they are both x to the fourth power

Angle:

what would you get if you added those two values?

gavinbacon2:

−11x^4

Angle:

awesome :)

gavinbacon2:

i have to answer quickly i have 11 minutes left on the test

Angle:

then combine \(5x^3\) with \(3x^3\) because they both have x to the third power

Angle:

we can't help on timed tests. it is against our anti-cheating policy

Angle:

I would like to point out that most of what you have is correct except for the negative in front of the 11 that we just got

gavinbacon2:

oh ok sorry

Angle:

no problem, I understand you're new here - feel free to ask any homework problem in the future :)

tiahnadomino:

(5+5x3-7x4) + (x+3x3-4x4-8)=43 5+5x ^3-7x^4+ x +3x^3-4x^4-8=43 -11x^4+8x^3+x-3=43 -11x^4+8x^3+x-3-43=43 -43 -11x^4+8x^3+x-46=0 Answer: NO REAL SOLUTIONS Tip: You can use mathpapa.com to help with problems like these